# Is Cognition Consistent with Perception? Assessing and Mitigating Multimodal Knowledge Conflicts in Document Understanding

### v2

@@ Abstract
Multimodal large language models (MLLMs) have shown impressive capabilities in document understanding, a rapidly growing research area with significant industrial demand. As a multimodal task, document understanding requires models to possess both perceptual and cognitive abilities. However, due to different types of annotation noise in training, current MLLMs often face conflicts between perception and cognition. Taking a document VQA task (cognition) as an example, an MLLM might generate answers that do not match the corresponding visual content identified by its OCR (perception). This conflict suggests that the MLLM might struggle to establish an intrinsic connection between the information it “sees” and what it “understands”. Such conflicts challenge the intuitive notion that cognition is consistent with perception, hindering the performance and explainability of MLLMs. In this paper, we define the conflicts between cognition and perception as Cognition and Perception (C&P) knowledge conflicts, a form of multimodal knowledge conflicts, and systematically assess them with a focus on document understanding. Our analysis reveals that even GPT-4o, a leading MLLM, achieves only 75.26% C&P consistency. To mitigate the C&P knowledge conflicts, we propose a novel method called Multimodal Knowledge Consistency Fine-tuning. Our method reduces C&P knowledge conflicts across all tested MLLMs and enhances their performance in both cognitive and perceptual tasks. All data we construct will be publicly available.
@@
Zirui Shao1thanks:    Equal contribution. , Feiyu Gao211footnotemark: 1, Zhaoqing Zhu211footnotemark: 1, Chuwei Luo2thanks:    Corresponding author., Hangdi Xing1, Zhi Yu122footnotemark: 2, Qi Zheng2, Ming Yan2 Jiajun Bu1 1Zhejiang University 2Alibaba Group shaozirui@zju.edu.cn
@@ 1 Introduction
In recent years, multimodal large language models (MLLMs) (gpt, 2023; Team et al., 2023; gpt, 2024; Chen et al., 2024; Bai et al., 2025; Ye et al., 2024; Li et al., 2024a) have witnessed rapid development and have demonstrated remarkable capabilities across a wide range of multimodal tasks (Antol et al., 2015; Mathew et al., 2021; Hossain et al., 2019). Of particular note is their application in document understanding (Cui et al., 2021; Xu et al., 2020, 2021; Huang et al., 2022; Luo et al., 2023), an area of high academic and industrial value, where significant progress has been made (Zhang et al., 2023a; Ye et al., 2023a, b; Luo et al., 2024; Wang et al., 2023; Hu et al., 2024).
As a multimodal task, document understanding requires models to accurately perceive visual content (perception, e.g., OCR) and then generate coherent responses (cognition, e.g., VQA) based on that perception. However, current MLLMs train perception and cognition using different sources of annotation (Bai et al., 2023; Hu et al., 2024). Perception typically relies on external OCR engines, while cognition often depends on human-annotated or LLM-generated data (Mathew et al., 2021; Van Landeghem et al., 2023). This discrepancy leads to different noise profiles, creating conflicts between perception and cognition. As shown in Figure 1, GPT-4o (gpt, 2024) recognizes the text in a certain region of an image as “Doral” but responds to a related VQA question with the text “Doraf”. This conflict suggests that GPT-4o struggles to establish a consistent connection between what it “sees” and what it “understands”. Statistical analysis further underscores this issue, as Figure 2 demonstrates that leading MLLMs like GPT-4o achieve only 75.26% consistency between perception and cognition (Section 3).
In this paper, we define intrinsic conflicts between cognitive knowledge and perceptual knowledge within MLLMs, which result in inconsistencies in responses related to cognition and perception, as Cognition and Perception (C&P) knowledge conflicts (Section 2.1). These conflicts undermine the explainability of MLLMs, as they challenge the intuitive notion that cognition is consistent with perception. Unlike previous research on multimodal knowledge conflicts (e.g., hallucination) (Zhai et al., 2024; Li et al., 2023; Guan et al., 2024; Liu et al., 2023), which focuses solely on conflicts within either cognition or perception, we highlight, for the first time, the conflicts that arise between the two.
@@ 2.1 The Definition of Cognition and Perception Knowledge Conflicts
In the training process of current MLLMs, annotations for perceptual tasks (e.g., OCR) and cognitive tasks (e.g., VQA) are often derived from different sources. For example, in the widely used DocVQA dataset (Mathew et al., 2021), OCR annotations are generated by commercial OCR solutions, while VQA annotations are crowd-sourced. Differences in annotation origins introduce discrepancies in noise and content, resulting in inconsistent bias that creates conflicts between cognitive and perceptual knowledge, referred to as Cognition and Perception (C&P) knowledge conflicts. Such conflicts manifest when $y_{C}$ and $y_{P}$ are inconsistent, i.e., $\delta(y_{C},y_{P})=0$. It is important to note that C&P knowledge conflicts do not consider whether $y_{C}=GT$ or $y_{P}=GT$. To quantify the severity of these conflicts, we introduce C&P consistency. Let $N$ denote the number of query pairs, with the C&P consistency calculated as follows:
[…]
In this paper, we focus on document understanding, where given a text $GT$ within $x_{I}$ bounded by $Box$, $x_{C}$ is a VQA query using $GT$ as the answer, and $x_{P}$ is an OCR query operating solely within $Box$. In practice, $Box$ may contain additional text besides $GT$. Consequently, C&P knowledge conflicts occur when $y_{P}$ does not fully contain $y_{C}$. The $\delta(y_{C},y_{P})$ can be specifically defined as follows:
@@
Furthermore, performance gaps may cause models to exhibit C&P inconsistency. For example, MLLMs may fail to comprehend VQA questions. Therefore, we introduce an auxiliary metric, called “Idealized C&P Consistency,” which evaluates inconsistencies only when both $ANLS(y_{C},GT)$ and $ANLS(y_{P},GT)$ are at least 0.5. The ANLS metric (Biten et al., 2019) is widely used in document understanding to measure text similarity on a scale from 0 to 1. Generally, cases with ANLS below 0.5 are considered complete failures of the model’s response to a query. By filtering out these poor cases caused by model performance, this metric provides additional insight into the C&P consistency under ideal conditions.
Figure 2: a: C&P knowledge conflicts in current MLLMs. “*” denotes the “SFT-baseline” (see Section 3). Additional qualitative results are provided in Section A.4 and Table 1. b: The distribution of conflict patterns, including character-level errors (P1), cognitive bias (P2), and limited cognitive ability (P3), with one illustrative example for each.
@@ 2.2 The Construction of Evaluation Samples
Given an image $x_{I}$ with its QA annotation $(Q,A)$, we assign $GT=A$ and $x_{C}=Q$. We construct $x_{P}$ using visual prompting (Wu et al., 2024b; Yang et al., 2023). $x_{P}$ is a simple question: “What is the text within the red box?” The corresponding image $x_{I}^{B}$ is obtained by drawing a red box in $x_{I}$ at the location of $Box$, denoted as $x_{I}^{B}=\text{VisP}(x_{I},Box)$, where $\text{VisP}(\cdot)$ represents the visual prompting process and $Box$ is the bounding box containing $GT$. In practice, responses for cognitive and perceptual tasks are obtained as $y_{C}=f(x_{C},x_{I})$ and $y_{P}=f(x_{P},x_{I}^{B})$, respectively.
@@
In particular, we consider five document understanding datasets to construct evaluation samples, which are categorized into three tasks: Document Question Answering (DocVQA (Mathew et al., 2021) and DUDE (Van Landeghem et al., 2023)), Document Information Extraction (DeepForm (Svetlichnaya, 2020) and FUNSD (Jaume et al., 2019)), and Chart Question Answering (ChartQA (Masry et al., 2022)). The evaluation samples are constructed from the test sets of these datasets. Section A.1 and A.2 provides additional details, including dataset descriptions, an example evaluation sample, and comprehensive statistics.
@@ 3 The Cognition and Perception Knowledge Conflicts in Current MLLMs
Two closed-source and three open-source MLLMs are evaluated. The closed-source models, GPT-4o (gpt, 2024) and Qwen-VL-Max (Bai et al., 2023, 2025), are well-regarded in the community. We evaluate these models using their publicly available APIs, disabling all randomness-inducing hyperparameters. Additionally, to ensure that MLLMs follow instructions, we carefully adjust the prompts based on the characteristics of each dataset. Details are provided in Section A.4.
The open-source models include InternVL2-2b (Chen et al., 2024), InternVL2-8b (Chen et al., 2024), and Qwen2.5-VL-7b (Bai et al., 2025), which differ in size and architecture. We perform the evaluation by disabling all randomness-inducing hyperparameters on an Nvidia A100 GPU. Furthermore, we observe that using the original weights for inference leads to issues with instruction following (see Section A.4 for details), and thus we construct SFT data using the training sets from all datasets following the procedure outlined in Section 2.2 to train baseline models for each MLLM, referred to as “SFT-baseline”. Training details are provided in Section 5.1.
Figure 2 (a) presents the evaluation results, with more qualitative results provided in Section A.4 and Table 1. Overall, C&P knowledge conflicts are common in current MLLMs, with inconsistencies observed in 12%–25% of cases. Furthermore, the severity of these conflicts appears comparable between open-source and closed-source models.
To gain a clearer picture of C&P knowledge conflicts, we randomly sample 10% of all inconsistent cases generated by InternVL2-2b and manually inspect them. We identify three main types of conflicts, as shown in Figure 2 (b). The majority of conflicts (41.3%, P1) stem from character-level errors when the model responds to either the OCR or VQA query. P2 (17.4%) arises from cognitive bias. Although the model “sees” the correct text (its OCR output is accurate), it prefers a linguistically more plausible answer (e.g., substituting “round tin packaging” with “round in packaging”). P1 and P2 together support our hypothesis that heterogeneous VQA and OCR annotations are a primary source of C&P knowledge conflicts. P1 reflects perception noise introduced by external OCR engine annotators, while P2 reflects semantic bias introduced by human or LLM annotators. P3 (34.8%) reveals a limitation in the model’s cognitive ability, where the VQA response is hallucinated despite an accurate OCR output. To focus on purer conflict conditions, we exclude P3 from the idealized C&P consistency (Section 2.1).
@@ 4 Multimodal Knowledge Consistency Fine-tuning
Section 3 demonstrates that even state-of-the-art MLLMs exhibit C&P knowledge conflicts. To resolve these conflicts, we propose Multimodal Knowledge Consistency Fine-tuning, illustrated in Figure 3, which comprises two components: C&P Link Tokens and the C&P Connector. As heterogeneous VQA and OCR annotations are a primary source of C&P conflicts (Section 3), this method aims to reinforce the connection between cognitive and perceptual tasks.
@@ 4.1 C&P Link Tokens
Previous research (Wu et al., 2024a) indicates that special tokens can effectively connect knowledge across different tasks. Therefore, we define a pair of C&P Link Tokens to connect cognitive and perceptual tasks, namely <CPLINK> and </CPLINK>, and add them to the original MLLM vocabulary. When the MLLM responds to a query using text extracted from an image, it encloses that text with the two C&P link tokens, for example, “<CPLINK>XXX</CPLINK>.” Given an image $x_{I}$ with QA annotation $(Q,A)$, the cognitive task’s query and response are $(x_{C},y_{C})$ and the perceptual task’s are $(x_{P},y_{P})$. According to Section 2.2, both $y_{C}$ and $y_{P}$ are texts derived from $x_{I}$, i.e. $A$. Therefore, the C&P link tokens can be applied to the responses of both tasks, denoted as $y_{C}^{\prime}$ and $y_{P}^{\prime}$, thereby strengthening their connection. Additionally, for guiding linked responses, we design $x_{C}^{\prime}$ to more explicitly prompt the model by adding a special instruction: “<CPLINK>XXX</CPLINK> indicates the OCR-derived answer.”
@@ 4.2 C&P Connector
DocVQA DUDE DeepForm FUNSD ChartQA Average
InternVL2-2b* 80.59 90.62 64.69 83.00 72.05 77.40 80.84 87.95 83.80 91.27 76.40 86.05
InternVL2-2b (Ours) 83.39 91.32 69.49 84.75 78.56 82.20 81.50 89.60 87.64 93.17 80.12 88.21
InternVL2-8b* 84.28 91.32 67.82 83.14 74.19 77.70 82.60 91.82 86.88 91.86 79.15 87.17
InternVL2-8b (Ours) 87.32 93.03 73.26 84.70 79.17 82.22 83.48 90.13 90.53 94.22 82.75 88.86
Qwen2.5-VL-7b* 93.79 96.44 79.30 91.87 75.20 85.36 84.80 90.38 92.06 95.37 85.03 91.88
Qwen2.5-VL-7b (Ours) 94.95 97.10 84.04 94.22 79.57 86.73 90.31 94.07 93.09 95.74 88.39 93.57

### v3

@@ Abstract
Multimodal large language models (MLLMs) have shown impressive capabilities in document understanding, a rapidly growing research area with significant industrial demand. As a multimodal task, document understanding requires models to possess both perceptual and cognitive abilities. However, due to different types of annotation noise in training, current MLLMs often face conflicts between perception and cognition. Taking a document VQA task (cognition) as an example, an MLLM might generate answers that do not match the corresponding visual content identified by its OCR (perception). This conflict suggests that the MLLM might struggle to establish an intrinsic connection between the information it “sees” and what it “understands”. Such conflicts challenge the intuitive notion that cognition is consistent with perception, hindering the performance and explainability of MLLMs. In this paper, we define the conflicts between cognition and perception as Cognition and Perception (C&P) knowledge conflicts, a form of multimodal knowledge conflicts, and systematically assess them with a focus on document understanding. Our analysis reveals that even GPT-4o, a leading MLLM, achieves only 75.26% C&P consistency. To mitigate the C&P knowledge conflicts, we propose a novel method called Multimodal Knowledge Consistency Fine-tuning. Our method reduces C&P knowledge conflicts across all tested MLLMs and enhances their performance in both cognitive and perceptual tasks.
@@
Zirui Shao1thanks:    Equal contribution., Feiyu Gao211footnotemark: 1, Zhaoqing Zhu211footnotemark: 1, Chuwei Luo2thanks:    Corresponding author., Hangdi Xing1, Zhi Yu1,322footnotemark: 2, Qi Zheng2, Ming Yan2, Jiajun Bu1 1 Zhejiang Key Laboratory of Accessible Perception and Intelligent Systems, Zhejiang University 2Alibaba Group, 3 Hangzhou High-Tech Zone (Binjiang) Institute of Blockchain and DataSecurity {shaozirui, yuzhirenzhe}@zju.edu.cn, feiyu.gfy@alibaba-inc.com {zzhaoqing.z, luochuwei}@gmail.com
@@ 1 Introduction
In recent years, multimodal large language models (MLLMs) (gpt4v; team2023gemini; gpt4o; internvl; bai2025qwen2; ye2024mplug; li2024llava) have witnessed rapid development and have demonstrated remarkable capabilities across a wide range of multimodal tasks (antol2015vqa; docvqa; hossain2019comprehensive). Of particular note is their application in document understanding (cui2021document; xu2020layoutlm; xu2021layoutlmv2; huang2022layoutlmv3; luo2023geolayoutlm), an area of high academic and industrial value, where significant progress has been made (zhang2023llavar; ye2023mplugdocowl; ye2023ureader; layoutllm; wang2023docllm; docowl15).
As a multimodal task, document understanding requires models to accurately perceive visual content (perception, e.g., OCR) and then generate coherent responses (cognition, e.g., VQA) based on that perception. However, current MLLMs train perception and cognition using different sources of annotation (qwen-vl; docowl15). Perception typically relies on external OCR engines, while cognition often depends on human-annotated or LLM-generated data (docvqa; dude). This discrepancy leads to different noise profiles, creating conflicts between perception and cognition. As shown in Figure 1, GPT-4o (gpt4o) recognizes the text in a certain region of an image as “Doral” but responds to a related VQA question with the text “Doraf”. This conflict suggests that GPT-4o struggles to establish a consistent connection between what it “sees” and what it “understands”. Statistical analysis further underscores this issue, as Figure 2 demonstrates that leading MLLMs like GPT-4o achieve only 75.26% consistency between perception and cognition (Section 3).
In this paper, we define intrinsic conflicts between cognitive knowledge and perceptual knowledge within MLLMs, which result in inconsistencies in responses related to cognition and perception, as Cognition and Perception (C&P) knowledge conflicts (Section 2.1). These conflicts undermine the explainability of MLLMs, as they challenge the intuitive notion that cognition is consistent with perception. Unlike previous research on multimodal knowledge conflicts (e.g., hallucination) (zhai2024halleswitch; li2023evaluating; guan2024hallusionbench; liu2023mitigating), which focuses solely on conflicts within either cognition or perception, we highlight, for the first time, the conflicts that arise between the two.
@@ 2.1 The Definition of Cognition and Perception Knowledge Conflicts
In the training process of current MLLMs, annotations for perceptual tasks (e.g., OCR) and cognitive tasks (e.g., VQA) are often derived from different sources. For example, in the widely used DocVQA dataset (docvqa), OCR annotations are generated by commercial OCR solutions, while VQA annotations are crowd-sourced. Differences in annotation origins introduce discrepancies in noise and content, resulting in inconsistent bias that creates conflicts between cognitive and perceptual knowledge, referred to as Cognition and Perception (C&P) knowledge conflicts. Such conflicts manifest when $y_{C}$ and $y_{P}$ are inconsistent, i.e., $\delta(y_{C},y_{P})=0$. It is important to note that C&P knowledge conflicts do not consider whether $y_{C}=GT$ or $y_{P}=GT$. To quantify the severity of these conflicts, we introduce C&P consistency. Let $N$ denote the number of query pairs, with the C&P consistency calculated as follows:
$\text{C\&P Consistency}=\frac{\sum_{i=1}^{N}\delta(y_{C_{i}},y_{P_{i}})}{N}.$ (1)
In this paper, we focus on document understanding and follow common practice (fu2024mmecomprehensiveevaluationbenchmark; chen2024pca) by using OCR as a representative perceptual task and VQA as a representative cognitive task. Specifically, given a text $GT$ within $x_{I}$ bounded by $Box$, $x_{C}$ is a VQA query using $GT$ as the answer, and $x_{P}$ is an OCR query operating solely within $Box$. In practice, $Box$ may contain additional text besides $GT$. Consequently, C&P knowledge conflicts occur when $y_{P}$ does not fully contain $y_{C}$. The $\delta(y_{C},y_{P})$ can be specifically defined as follows:
@@
Furthermore, performance gaps may cause models to exhibit C&P inconsistency. For example, MLLMs may fail to comprehend VQA questions. Therefore, we introduce an auxiliary metric, called “Idealized C&P Consistency,” which evaluates inconsistencies only when both $ANLS(y_{C},GT)$ and $ANLS(y_{P},GT)$ are at least 0.5. The ANLS metric (biten2019scene) is widely used in document understanding to measure text similarity on a scale from 0 to 1. Generally, cases with ANLS below 0.5 are considered complete failures of the model’s response to a query. By filtering out these poor cases caused by model performance, this metric provides additional insight into the C&P consistency under ideal conditions.
Figure 2: a: C&P knowledge conflicts in current MLLMs. “*” denotes the “SFT-baseline” (see Section 3). Additional quantitative results are provided in Section A.4 and Table 1. b: Results of the synthetic noise experiment, with additional details provided in Section A.5. c: The distribution of conflict patterns, including character-level errors (P1), cognitive bias (P2), and limited cognitive ability (P3), with one illustrative example for each.
@@ 2.2 The Construction of Evaluation Samples
Given an image $x_{I}$ with its QA annotation $(Q,A)$, we assign $GT=A$ and $x_{C}=Q$. We construct $x_{P}$ using visual prompting (wu2024visual; yang2023set). $x_{P}$ is a simple question: “What is the text within the red box?” The corresponding image $x_{I}^{B}$ is obtained by drawing a red box in $x_{I}$ at the location of $Box$, denoted as $x_{I}^{B}=\text{VisP}(x_{I},Box)$, where $\text{VisP}(\cdot)$ represents the visual prompting process and $Box$ is the bounding box containing $GT$. In practice, responses for cognitive and perceptual tasks are obtained as $y_{C}=f(x_{C},x_{I})$ and $y_{P}=f(x_{P},x_{I}^{B})$, respectively.
@@
In particular, we consider five document understanding datasets to construct evaluation samples, which are categorized into three tasks: Document Question Answering (DocVQA (docvqa) and DUDE (dude)), Document Information Extraction (DeepForm (deepform) and FUNSD (funsd)), and Chart Question Answering (ChartQA (chartqa)). The evaluation samples are constructed from the test sets of these datasets. Section A.1 and A.2 provides additional details, including dataset descriptions, an example evaluation sample, and comprehensive statistics.
@@ 3 The Cognition and Perception Knowledge Conflicts in Current MLLMs
Two closed-source and three open-source MLLMs are evaluated. The closed-source models, GPT-4o (gpt4o) and Qwen-VL-Max (qwen-vl; bai2025qwen2), are well-regarded in the community. We evaluate these models using their publicly available APIs, disabling all randomness-inducing hyperparameters. Additionally, to ensure that MLLMs follow instructions, we carefully adjust the prompts based on the characteristics of each dataset. Details are provided in Section A.4.
The open-source models include InternVL2-2b (internvl), InternVL2-8b (internvl), and Qwen2.5-VL-7b (bai2025qwen2), which differ in size and architecture. We perform the evaluation by disabling all randomness-inducing hyperparameters on an Nvidia A100 GPU. Furthermore, we observe that using the original weights for inference leads to issues with instruction following (see Section A.4 for details), and thus we construct SFT data using the training sets from all datasets following the procedure outlined in Section 2.2 to train baseline models for each MLLM, referred to as “SFT-baseline”. Training details are provided in Section 5.1.
Figure 2 (a) presents the evaluation results, with more quantitative results provided in Section A.4 and Table 1. Overall, C&P knowledge conflicts are common in current MLLMs, with inconsistencies observed in 12%–25% of cases. Furthermore, the severity of these conflicts appears comparable between open-source and closed-source models.
To further investigate the potential cause of C&P knowledge conflicts, we train InternVL2-2b with varying levels of synthetic noise (OCR: shape mix-ups, missing or extra letters; VQA: typos, omitted details). Synthetic noise is injected into 5%, 10%, and 20% of the training data. Figure 2 (b) shows the results, with additional quantitative analysis provided in Section A.5. Overall, as the level of noise increases, the C&P consistency declines.
We also randomly sample 10% of all inconsistent cases generated by InternVL2-2b and manually inspect them. Three main types of conflicts are identified, as shown in Figure 2 (c). The majority of conflicts (41.3%, P1) stem from character-level errors when the model responds to either the OCR or VQA query. P2 (17.4%) arises from cognitive bias. Although the model “sees” the correct text (its OCR output is accurate), it prefers a linguistically more plausible answer (e.g., substituting “round tin packaging” with “round in packaging”). The synthetic noise experiment, together with P1 and P2, supports our hypothesis that heterogeneous VQA and OCR annotations are a primary source of C&P knowledge conflicts. Specifically, P1 reflects perception noise introduced by external OCR engine annotators, while P2 reflects semantic bias introduced by human or LLM annotators. P3 (34.8%) reveals a limitation in the model’s cognitive ability, where the VQA response is hallucinated despite an accurate OCR output. To focus on purer conflict conditions, we exclude P3 from the idealized C&P consistency (Section 2.1).
@@ 4 Multimodal Knowledge Consistency Fine-tuning
Section 3 demonstrates that even state-of-the-art MLLMs exhibit C&P knowledge conflicts. To resolve these conflicts, we propose Multimodal Knowledge Consistency Fine-tuning, illustrated in Figure 3, which comprises two components: C&P Link Tokens and the C&P Connector. As heterogeneous VQA and OCR annotations are a primary source of C&P conflicts (Section 3), this method aims to reinforce the connection between cognitive and perceptual tasks, thereby mitigating C&P knowledge conflicts.
DocVQA DUDE DeepForm FUNSD ChartQA Average
InternVL2-2b*
80.59
90.62
64.69
83.00
72.05
77.40
80.84
87.95
83.80
91.27
76.40
86.05
InternVL2-2b (Ours)
83.39
91.32
69.49
84.75
78.56
82.20
81.50
89.60
87.64
93.17
80.12
88.21
InternVL2-8b*
84.28
91.32
67.82
83.14
74.19
77.70
82.60
91.82
86.88
91.86
79.15
87.17
InternVL2-8b (Ours)
87.32
93.03
73.26
84.70
79.17
82.22
83.48
90.13
90.53
94.22
82.75
88.86
Qwen2.5-VL-7b*
93.79
96.44
79.30
91.87
75.20
85.36
84.80
90.38
92.06
95.37
85.03
91.88
Qwen2.5-VL-7b (Ours)
94.95
97.10
84.04
94.22
79.57
86.73
90.31
94.07
93.09
95.74
88.39
93.57
@@ 4.1 C&P Link Tokens
Previous research (wu2024visionllm) indicates that special tokens can effectively connect knowledge across different tasks. Therefore, we define a pair of C&P Link Tokens to connect cognitive and perceptual tasks, namely <CPLINK> and </CPLINK>, and add them to the original MLLM vocabulary. When the MLLM responds to a query using text extracted from an image, it encloses that text with the two C&P link tokens, for example, “<CPLINK>XXX</CPLINK>.” Given an image $x_{I}$ with QA annotation $(Q,A)$, the cognitive task’s query and response are $(x_{C},y_{C})$ and the perceptual task’s are $(x_{P},y_{P})$. According to Section 2.2, both $y_{C}$ and $y_{P}$ are texts derived from $x_{I}$, i.e. $A$. Therefore, the C&P link tokens can be applied to the responses of both tasks, denoted as $y_{C}^{\prime}$ and $y_{P}^{\prime}$, thereby strengthening their connection. Additionally, for guiding linked responses, we design $x_{C}^{\prime}$ to more explicitly prompt the model by adding a special instruction: “<CPLINK>XXX</CPLINK> indicates the OCR-derived answer.”
@@ 4.2 C&P Connector
$\mathcal{X}=\{$ $(x^{\prime}_{C_{i}},y^{\prime}_{C_{i}}),(x_{P_{i}},y^{\prime}_{P_{i}}),$ (5)
$(x_{Conn_{i}}^{+},y_{Conn_{i}}^{+}),(x_{Conn_{i}}^{-},y_{Conn_{i}}^{-})\}_{i=0}^{N}.$